OpenStudy (anonymous):

how do you solve polynomial equations by factoring when there is a number without a variable at the end such as: y^2+ 12y+ 27= 0 ?

OpenStudy (amistre64):

have you tried subtracting it to the other side?

OpenStudy (amistre64):

the trick is to add 0

OpenStudy (amistre64):

but a 0 that is useful

OpenStudy (amistre64):

see that number next to the "y" ... the 12? halve it and square the results

OpenStudy (amistre64):

y^2+ 12y+ 27= 0 ; 12/2 = 6, 6^2 = 36 y^2 +12y +36 -36 +27 = 0 (y^2 +12y +36) -36 +27 = 0 (y+6)^2 -9 = 0

OpenStudy (anonymous):

well, this is as far as ive gotten: y (y+12) + 27 = 0 i was told to factor out and the only thing i saw was the "Y" but the 27 didnt have a "Y" so i just left it out.

OpenStudy (amistre64):

you did fine, but that way doesnt really get us to something useful; so we have to use other tricks that math provides

OpenStudy (diyadiya):

isn't it (y+3)(y+9) ??

OpenStudy (anonymous):

i think what we're trying to find is what numbers could "Y" be to cause the equation to equal 0.

OpenStudy (anonymous):

it may be but im not sure how to get there. :/

OpenStudy (amistre64):

(y+6)^2 -9 = 0 (y+6)^2 = 9 ; now we sqrt both sides y+6 = +- sqrt(9) y+6 = +- 3 y = -6 -3 , or y = -6 + 3

OpenStudy (amistre64):

your trying to think linearly; but this isnt a line equation; its a curve, a parabola to be exact

OpenStudy (diyadiya):

:D

OpenStudy (amistre64):

try to read up on a method called "completeing the square"